# Are LLMs Really Not Knowledgable? Mining the Submerged Knowledge in LLMs’ Memory

### v1

@@ Abstract
Large language models (LLMs) have shown promise as potential knowledge bases, yet they often struggle with question-answering tasks and are prone to hallucinations. While previous research attributes these issues to knowledge gaps in the model’s parameters, our investigation reveals a different phenomenon: LLMs often retain correct knowledge even when generating incorrect answers. Through analysis of model’s internal representations, we find that correct answers frequently appear among high-probability tokens despite not being selected as final outputs. Based on this observation, we introduce Hits@$k$, a new metric to assess knowledge retention independent of expression accuracy. Our extensive experiments demonstrate that LLMs store significantly more knowledge than their QA performance suggests. Building on these findings, we develop SkipUnsure, a method to improve answer accuracy by leveraging detected but unexpressed knowledge. Experiments on both open-domain and specific-domain datasets show consistent improvements, with accuracy gains of up to 11.8% on DBPedia and 6.3% on IMDB, without requiring model retraining.
@@ 1 Introduction
Based on these insights, we develop SkipUnsure, an approach that improves answer accuracy by effectively utilizing knowledge present in model representations. Unlike existing methods requiring substantial computational resources or architectural changes, our approach identifies and filters uninformative responses while recovering high-probability knowledge from internal distributions. Experiments demonstrate consistent improvements across diverse scenarios, achieving accuracy gains of 11.8% on DBPedia and 6.3% on IMDB. These improvements are particularly pronounced for head entities where models show strong knowledge retention but poor expression.
This work makes several key contributions:
  • Identification and analysis of the systematic gap between knowledge storage and expression in LLMs;
  • Introduction of Hits@$k$ for quantifying model knowledge retention;
  • Comprehensive examination of factors influencing knowledge storage and expression patterns;
  • Development of SkipUnsure, demonstrating practical benefits of leveraging stored but unexpressed knowledge.
@@ 2.1 Knowledge Storage and Expression
Our investigation reveals that model failures may stem from expression issues rather than knowledge gaps. Through systematic analysis of model outputs, we find that LLMs often retain correct information in their parameters even when generating incorrect answers. As shown in Figure 1, when asked about Washington state’s capital, while the model outputs “Seattle”, it assigns a high probability score to the correct answer “Olympia”. Such cases indicate the need for a deeper understanding of how knowledge is stored and expressed in these models.
@@ 2.3 The Hits@$k$ Metric
Experimental results in Figure 4 demonstrate the effectiveness of this metric in revealing stored knowledge. Using LLaMA3-8b on DBpedia, while Hits@1 is only 17.2%, Hits@5 reaches 57.9%, indicating substantially more stored knowledge than suggested by traditional metrics. This pattern holds across different domains and data types, suggesting a fundamental characteristic of how LLMs store and access information. These findings motivate a deeper examination of factors affecting knowledge storage and expression, which we explore in Section 3.
@@ 3.2 Models and Implementation
We conduct experiments using the following LLMs: LLaMA2-13b, LLaMA2-70b, LLaMA3-8b, LLaMA3-70b, LLaMA3.1-8b, Qwen2-1.5b, Qwen2-7b, Qwen2-72b. These models represent different architectural choices and parameter scales, ranging from 1.5B to 70B parameters. To minimize randomness in model outputs, we use greedy decoding with temperature set to 0.0 across all experiments.
@@ 4.1.1 Hits@$k$ performance of different models
Figure 2 shows the results on dataset DBPedia-head, demonstrating this finding. As the number of parameters increases, LLMs exhibit improved accuracy across a range of tasks, including QA tasks. This is due to the greater representational power of larger models, allowing them to capture more nuanced and complex language patterns. However, among the three datasets used for testing, the Hits@$k$ results for the LLaMA2-13b and LLaMA2-70b models are similar, as are the Hits@$k$ results for the LLaMA3-8b and LLaMA3-70b models. As shown in Figure 3, the rankings of LLMs based on Accuracy and Hits@$k$ differ significantly. This indicates that increasing the model size does not necessarily lead to richer or more comprehensive memory in LLMs.
Figure 2: The Hits@$k$ scores of different large language models on the DBPedia-head dataset.
(a) LLMs sorted by Accuracy
(b) LLMs sorted by Hits@$k$
Figure 3: The ranking of LLMs based on Accuracy and Hits@$k$ on DBPedia-Head.
@@ 4.1.2 Analysis of the influence of $k$ value selection
Figure 4 presents the Hits@$k$ scores for various $k$ values. A lower Hits@1 score suggests that the model struggles to provide the correct answer directly in the QA task. However, as the $k$ value increases, the score improves, indicating that the model retains relevant knowledge. Our experimental results indicate that, when $k=50$, the Hits@$k$ for the head, torso, and tail sections exceeds 80%. Despite the extensive vocabulary of the LLaMA3 model (approximately 128,000 tokens), the correct answer is frequently located within a relatively small number of tokens at the beginning. This suggests that the model has the potential to provide correct answers in most cases, even if an incorrect answer is initially generated. We observe a significant difference between scores at $k=1$ and $k=5$, indicating that utilizing tokens with higher probabilities can yield more reliable answers.
We show the cumulative distribution of the ranks of Hits@$k$ score in the QA task in Figure 5. We observe that the difference in popularity has a smaller impact on Hits@$k$ for the DBPedia dataset compared to IMDB. This suggests that the domain of datasets influences the sensitivity to popularity. Generally, memory performance on open-domain datasets is less sensitive to variations in popularity.
(a) DBPedia on LLaMA3-8b
(b) IMDB on LLaMA3-8b
@@ 4.2.1 Comparison of Open Domain and Specific Domain
As shown in Table 1, The Hits@$k$ results of data sets in different domains are different. DBLP is an open-domain dataset, while IMDB and Goodreads are domain-specific datasets. The experimental results demonstrate that the Hits@$k$ for the open-domain dataset is higher than that for the domain-specific datasets.
We show the cumulative distribution of the ranks of Hits@$k$ score in the QA task in Figure 5. We observe that the difference in popularity has a smaller impact on Hits@$k$ for the DBPedia dataset compared to IMDB. This suggests that the domain of the dataset influences the sensitivity to popularity. Generally, memory performance on open-domain datasets is less sensitive to variations in popularity.
@@ 4.4 Uninformative Response Impact
To reduce the likelihood of the model providing incorrect answers, prompts in QA tasks often include an “unsure” option, allowing the model to respond with “unsure” when uncertain about the correct answer. This approach helps minimize the risk of hallucinations when the model encounters unfamiliar information. Figure 7 shows the distribution ratios of three response types under the LLaMA3-8b model: uninformative, correct, and wrong. We found that uninformative responses have a greater impact on the model’s performance. Our experiments revealed that when some models answered “unsure”, they still retained relevant knowledge in memory. This suggests that the model may respond with “unsure” even when relevant memory exists. We show an example of this situation in Figure 6.
@@
In summary, we classify the cases mentioned above as uninformative responses, with Figure 6 showing the proportion of such responses across different datasets. We show two different types of uninformative responses in Figure 6, the correct answers appear in the tokens corresponding to the second-highest or third-highest logits. In Questions 1 and 2, the model’s final answer is “unsure,” yet the correct answer, or a related subword, appears as the second most probable token in the logits. In Question 3, the model’s final answer is a null character, which is considered incorrect. However, the correct answer is found in the token with the third-highest probability. This indicates that, while the model possesses relevant memory, it fails to output the correct answer.
As shown in Figure 7, in the DBPedia dataset, experimental results show that more than half of the responses in the Head, Torso, and Tail sub-datasets are uninformative. In the domain-specific IMDB dataset, the high proportion of uninformative responses also significantly impacts the model’s accuracy in QA tasks. This highlights the significant impact of uninformative responses on the final results in both open-domain and domain-specific datasets. Moreover, as dataset popularity decreases, the proportion of uninformative responses increases, which emerges as a key factor contributing to the decline in accuracy in QA tasks.
@@
5 Method: SkipUnsure
Algorithm 1 SkipUnsure
1:Inputs: Output’s Token list $L$ (sorted by logits rank)
2:$i=0$
3:while $L[i]$ is uninformative do
4:     Delete $L[i]$
5:     $i=i+1$
6:end while
7:$\text{Prompt}_{new}\leftarrow\text{Prompt}_{old}+L[i]$
8:$\text{Output}_{new}\leftarrow LLM(\text{Prompt}_{new})$
5.1 Methodology
Our analysis in Section 4 demonstrates that LLMs often retain correct knowledge in their parameters even when generating incorrect answers. Particularly, we observe that models frequently output “unsure” or uninformative responses while assigning high probabilities to correct answers in their internal distributions. This finding suggests an opportunity to improve model performance by better utilizing these internal representations.
Motivated by this observation, we propose SkipUnsure, a method to effectively recover stored knowledge from model’s internal representations. Given a question $q$, let $P(t\mid q)$ denote the probability distribution over the model’s vocabulary V, and $T_{k}=\{t_{1},...,t_{k}\}$ be the top-$k$ tokens ranked by their logit scores.
Our method first identifies uninformative tokens through a filtering process. We define a token $t$ as uninformative if it starts with “uns”, represents an empty string, contains less than three characters, or consists only of stop words. This filtering strategy effectively removes low-confidence and meaningless responses while retaining potentially valuable information.
For knowledge recovery, we examine the remaining tokens in $T_{k}$ after filtering out the uninformative set $U$. The final answer $a^{*}$ is selected as the highest-probability token from the valid candidates:
$a^{*}=\arg\max_{t\in T_{k}\setminus U}P(t\mid q)$ (2)
Token $a^{*}$ will be appended to the end of the original prompt and re-entered into the model to generate a new response. This approach allows us to bypass uninformative responses and leverage the model’s stored knowledge that might otherwise be obscured.
The effectiveness of SkipUnsure stems from two key insights from our analysis. First, correct answers often appear in $T_{k}$ with high probability scores even when the model outputs uninformative responses. Second, uninformative tokens can be reliably identified through our filtering criteria. The method requires no model retraining or architectural modifications, making it easy to integrate with existing LLM deployments. Our algorithm is shown in Algorithm 1.
5.2 Experiments result
Following the experimental evaluation settings outlined in the previous article, we conducted evaluations on multiple base models, including those of varying sizes and both recent and older release dates. The prompts used are also consistent with those described in the previous experiments. As shown in Table 3 and Table 3, our experiments demonstrate that our method achieves performance improvements across datasets with varying levels of popularity. The improvement is particularly pronounced in datasets with higher popularity, suggesting that the model has more extensive memory storage in datasets with greater popularity. Using the LLaMA3-8b model, our method improved accuracy on the DBPedia dataset by 3.8% in the head, 3.2% in the torso, and 2.5% in the tail sections. With the LLaMA3-70b model, the improvements were 11.8%, 9.4%, and 6.7% in the respective sections. On the IMDB dataset, with the LLaMA3-70b model, the improvements were 6.3%, 5.4%, and 0.6% in the respective sections. Additionally, our experimental results indicate that uninformative responses impact the model’s memory expression to some extent. By addressing these uninformative responses, the model’s ability to express memory can be enhanced.
DBPedia w/o SkipUnsure SkipUnsure
@@ 5.2 Experiments result
LLaMA2-13b 8.4 4.0 3.2 13.5 $5.1$ 8.9 $4.9$ 6.9 $3.7$
LLaMA2-70b 16.0 12.3 8.1 18.3 $2.3$ 14.0 $1.7$ 9.4 $1.3$
LLaMA3-8b 9.8 7.3 5.1 13.6 $3.8$ 10.5 $3.2$ 7.6 $2.5$
LLaMA3-70b 11.2 8.7 6.0 23.0 $11.8$ 18.1 $9.4$ 12.7 $6.7$
LLaMA3.1-8b 8.1 5.3 3.7 15.6 $7.5$ 10.1 $4.8$ 7.4 $3.7$
Qwen2-1.5b 1.2 0.7 0.6 2.4 $1.2$ 1.5 $0.8$ 1.4 $0.8$
Qwen2-7b 3.9 2.5 1.3 9.3 $5.4$ 7.5 $5.0$ 4.8 $3.5$
Qwen2-72b 17.3 12.1 9.0 20.1 $2.8$ 14.3 $2.2$ 10.2 $1.2$
Table 2: Performance results of SkipUnsure on different base models of dataset DBPedia.
IMDB w/o SkipUnsure SkipUnsure
@@
LLaMA2-13b 15.7 11.1 0.0 21.9 $6.2$ 17.6 $6.5$ 2.0 $2.0$
LLaMA2-70b 25.2 23.5 4.3 25.2 23.5 4.5 $0.2$
[…]
LLaMA3-70b 19.1 18.6 4.3 25.4 $6.3$ 24.0 $5.4$ 4.9 $0.6$

### v2

# Are LLMs Really Not Knowledgeable? Mining the Submerged Knowledge in LLMs’ Memory

@@ Abstract
Large language models (LLMs) have shown promise as parametric knowledge bases, but often underperform on question answering (QA) tasks due to hallucinations and uncertainty. While prior work attributes these failures to knowledge gaps in the model’s parameters, we uncover a complementary phenomenon: LLMs frequently retain correct knowledge even when generating incorrect or “unsure” answers. By analyzing the token-level output distributions, we find that correct answers often appear among high-probability candidates, despite not being selected. Motivated by this, we propose Hits@$k$, a novel metric to evaluate latent knowledge retention independent of answer surface form. Our experiments reveal that LLMs possess significantly more factual knowledge than is reflected by standard QA accuracy. Building on these insights, we further examine the prevailing few-shot QA paradigm. We find that prompting strategies which allow “unsure” outputs can inadvertently suppress correct answers by discouraging low-confidence generation. We design a set of quantitative experiments to measure this suppression effect, offering practical guidance for future prompt and decoding design in knowledge-intensive tasks.
@@ 1 Introduction
Building on these insights, we further examine the widely adopted few-shot QA paradigm. We observe that prompting strategies which permit “unsure” responses may inadvertently suppress the generation of low-confidence answers, thereby inhibiting the expression of correct knowledge. In particular, we find that when the model outputs “unsure”, tokens ranked highly in the logit distribution—though not selected as the top-1 prediction—often contain the correct answer. To quantify this effect, we design a set of experiments where “unsure”-related tokens are filtered out during decoding. Under this strategy, a subset of previously masked correct answers can be successfully recovered. These findings highlight a potential trade-off between cautious generation and knowledge expressiveness, and provide actionable guidance for future prompt design and decoding strategies in knowledge-intensive tasks.
This work makes the following key contributions:
  • We identify and analyze a systematic gap between knowledge storage and expression in large language models (LLMs).
  • We propose Hits@$k$, a novel metric for quantifying latent knowledge retention independent of output accuracy.
  • We conduct a comprehensive analysis of the factors that influence the alignment between stored knowledge and generated answers.
  • We quantitatively demonstrate that existing prompting paradigms can inadvertently suppress correct answers by enabling “unsure” responses, revealing a memory-masking effect.
@@ 2.1 Knowledge Storage and Expression
Our investigation reveals that model failures may stem from expression issues rather than knowledge gaps. Through systematic analysis of model outputs, we find that LLMs often retain correct information in their parameters even when generating incorrect answers. As shown in Figure˜1, when asked about Washington state’s capital, while the model outputs “Seattle”, it assigns a high probability score to the correct answer “Olympia”. Such cases indicate the need for a deeper understanding of how knowledge is stored and expressed in these models.
@@ 2.3 The Hits@$k$ Metric
Experimental results in Figure˜4 demonstrate the effectiveness of this metric in revealing stored knowledge. Using LLaMA3-8b on DBpedia, while Hits@1 is only 17.2%, Hits@5 reaches 57.9%, indicating substantially more stored knowledge than suggested by traditional metrics. This pattern holds across different domains and data types, suggesting a fundamental characteristic of how LLMs store and access information. These findings motivate a deeper examination of factors affecting knowledge storage and expression, which we explore in Section˜3.
@@ 3.2 Models and Implementation
We conduct experiments using the following LLMs: LLaMA2-13b, LLaMA2-70b, LLaMA3-8b, LLaMA3-70b, LLaMA3.1-8b, Qwen2-1.5b, Qwen2-7b, Qwen2-72b, Mistral-7B-Instruct-v0.3 (abbreviated as Mistral-7b). These models represent different architectural choices and parameter scales, ranging from 1.5B to 70B parameters. To minimize randomness in model outputs, we use greedy decoding with temperature set to 0.0 across all experiments.
@@ 4.1.1 Hits@$k$ performance of different models
Figure˜2 shows the results on dataset DBPedia-head, demonstrating this finding. As the number of parameters increases, LLMs exhibit improved accuracy across a range of tasks, including QA tasks. This is due to the greater representational power of larger models, allowing them to capture more nuanced and complex language patterns. However, among the three datasets used for testing, the Hits@$k$ results for the LLaMA2-13b and LLaMA2-70b models are similar, as are the Hits@$k$ results for the LLaMA3-8b and LLaMA3-70b models. As shown in Figure˜3, the rankings of LLMs based on Accuracy and Hits@$k$ differ significantly. This indicates that increasing the model size does not necessarily lead to richer or more comprehensive memory in LLMs.
Figure 2: The Hits@$k$ scores of different large language models on the DBPedia-Head dataset when $k=100$.
((a))
((b))
Figure 3: The ranking of LLMs based on Accuracy and Hits@$k$ on DBPedia-Head when $k=100$.
@@
Justification for Hits@k as a Knowledge Metric.
We posit that Hits@k captures genuine latent knowledge rather than surface-level token co-occurrence. This validity is supported by three key observations: (1) the systematic nature of the storage-expression gap across diverse domains Figure˜2; (2) the distinct model rankings yielded by Hits@k compared to standard accuracy, indicating it measures a fundamental internal property Figure˜3, which empirically proves these tokens represent accessible, usable knowledge suppressed by decoding dynamics.
@@ 4.1.2 Analysis of the influence of $k$ value selection
Figure˜4 presents the Hits@$k$ scores for various $k$ values. A lower Hits@1 score suggests that the model struggles to provide the correct answer directly in the QA task. However, as the $k$ value increases, the score improves, indicating that the model retains relevant knowledge. Our experimental results indicate that, when $k=50$, the Hits@$k$ for the head, torso, and tail sections exceeds 80%. Despite the extensive vocabulary of the LLaMA3 model (approximately 128,000 tokens), the correct answer is frequently located within a relatively small number of tokens at the beginning. This suggests that the model has the potential to provide correct answers in most cases, even if an incorrect answer is initially generated. We observe a significant difference between scores at $k=1$ and $k=5$, indicating that utilizing tokens with higher probabilities can yield more reliable answers.
We show the cumulative distribution of the ranks of Hits@$k$ score in the QA task in Figure˜5. We observe that the difference in popularity has a smaller impact on Hits@$k$ for the DBPedia dataset compared to IMDB. This suggests that the domain of datasets influences the sensitivity to popularity. Generally, memory performance on open-domain datasets is less sensitive to variations in popularity.
((a))
((b))
@@ 4.2.1 Comparison of Open Domain and Specific Domain
As shown in Table˜1, The Hits@$k$ results of data sets in different domains are different. DBLP is an open-domain dataset, while IMDB and Goodreads are domain-specific datasets. The experimental results demonstrate that the Hits@$k$ for the open-domain dataset is higher than that for the domain-specific datasets.
We show the cumulative distribution of the ranks of Hits@$k$ score in the QA task in Figure˜5. We observe that the difference in popularity has a smaller impact on Hits@$k$ for the DBPedia dataset compared to IMDB. This suggests that the domain of the dataset influences the sensitivity to popularity. Generally, memory performance on open-domain datasets is less sensitive to variations in popularity.
@@ 4.4 Uninformative Response Impact
To reduce the likelihood of the model providing incorrect answers, prompts in QA tasks often include an “unsure” option, allowing the model to respond with “unsure” when uncertain about the correct answer. This approach helps minimize the risk of hallucinations when the model encounters unfamiliar information. Figure˜7 shows the distribution ratios of three response types under the LLaMA3-8b model: uninformative, correct, and wrong. We found that uninformative responses have a greater impact on the model’s performance. Our experiments revealed that when some models answered “unsure”, they still retained relevant knowledge in memory. This suggests that the model may respond with “unsure” even when relevant memory exists. We show an example of this situation in Figure˜6.
@@
In summary, we classify the cases mentioned above as uninformative responses, with Figure˜6 showing the proportion of such responses across different datasets. We show two different types of uninformative responses in Figure˜6, the correct answers appear in the tokens corresponding to the second-highest or third-highest logits. In Questions 1 and 2, the model’s final answer is “unsure,” yet the correct answer, or a related subword, appears as the second most probable token in the logits. In Question 3, the model’s final answer is a null character, which is considered incorrect. However, the correct answer is found in the token with the third-highest probability. This indicates that, while the model possesses relevant memory, it fails to output the correct answer.
As shown in Figure˜7, in the DBPedia dataset, experimental results show that more than half of the responses in the Head, Torso, and Tail sub-datasets are uninformative. In the domain-specific IMDB dataset, the high proportion of uninformative responses also significantly impacts the model’s accuracy in QA tasks. This highlights the significant impact of uninformative responses on the final results in both open-domain and domain-specific datasets. Moreover, as dataset popularity decreases, the proportion of uninformative responses increases, which emerges as a key factor contributing to the decline in accuracy in QA tasks.
@@
5 Revisiting “Unsure” Responses in Knowledge-Based QA
Table 2: Answer recovery rates from “unsure” responses on DBPedia (left) and IMDB (right) datasets. Filtering uninformative tokens reveals a substantial portion of correct answers masked during initial decoding.
DBPedia Greedy decoding Decoding with Unsure filter
@@ 5 Revisiting “Unsure” Responses in Knowledge-Based QA
LLaMA2-13b 8.4 4.0 3.2 13.5 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}5.1$ 8.9 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}4.9$ 6.9 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}3.7$
LLaMA2-70b 16.0 12.3 8.1 18.3 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}2.3$ 14.0 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}1.7$ 9.4 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}1.3$
LLaMA3-8b 9.8 7.3 5.1 13.6 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}3.8$ 10.5 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}3.2$ 7.6 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}2.5$
LLaMA3-70b 11.2 8.7 6.0 23.0 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}11.8$ 18.1 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}9.4$ 12.7 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}6.7$
LLaMA3.1-8b 8.1 5.3 3.7 15.6 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}7.5$ 10.1 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}4.8$ 7.4 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}3.7$
Qwen2-1.5b 1.2 0.7 0.6 2.4 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}1.2$ 1.5 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.8$ 1.4 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.8$
Qwen2-7b 3.9 2.5 1.3 9.3 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}5.4$ 7.5 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}5.0$ 4.8 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}3.5$
Qwen2-72b 17.3 12.1 9.0 20.1 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}2.8$ 14.3 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}2.2$ 10.2 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}1.2$
Mistral-7b 16.5 11.0 7.5 16.7 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.2$ 11.3 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.3$ 7.9 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.4$
IMDB Greedy decoding Decoding with Unsure filter
@@
LLaMA2-13b 15.7 11.1 0.0 21.9 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}6.2$ 17.6 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}6.5$ 2.0 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}2.0$
LLaMA2-70b 25.2 23.5 4.3 25.2 23.5 4.5 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.2$
LLaMA3-8b 20.7 16.4 2.2 21.3 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.6$ 17.1 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.7$ 2.7 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.5$
LLaMA3-70b 19.1 18.6 4.3 25.4 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}6.3$ 24.0 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}5.4$ 4.9 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.6$
LLaMA3.1-8b 18.3 13.4 2.0 18.8 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.5$ 14.5 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}1.1$ 2.3 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.3$
Qwen2-1.5b 2.0 0.9 0.5 2.5 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.5$ 1.1 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.2$ 0.5
Qwen2-7b 7.4 2.9 0.3 11.7 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}4.3$ 4.5 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}1.6$ 0.4 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.1$
Qwen2-72b 19.1 16.2 1.1 20.3 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}2.2$ 18.1 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}1.9$ 1.2 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.1$
Mistral-7b 20.5 15.1 1.3 20.5 15.2 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.1$ 1.4 $\uparrow$$\raisebox{0.5pt}{$\uparrow$}0.1$
5.1 Observation
In many knowledge-based question answering (KBQA) benchmarks, models are permitted to respond with “unsure” when they lack confidence in producing a correct answer. This strategy is commonly employed to reduce the risk of hallucinations or factually incorrect outputs. As illustrated in Figure˜7, a non-negligible portion of model predictions fall into this uninformative category.
However, this raises an important question: when a model generates “unsure”, does it truly lack the relevant knowledge, or is the correct answer being suppressed due to low decoding confidence? To investigate this, we analyze the token-level logit distributions during generation. Surprisingly, we observe that in a significant number of “unsure” cases, the correct answer still appears among the top-$k$ (e.g., $k=2$ or $3$) candidates by logit rank, even though the model ultimately selects the “unsure” token as the output.
5.2 Quantitative analysis
The observations above suggest that LLMs may retain correct knowledge even when they abstain from answering. To systematically measure this phenomenon, we design a set of controlled experiments to quantify the extent to which correct answers are recoverable from “unsure” outputs.
Our goal is to evaluate the latent presence of correct answers in the model’s internal distributions, independent of what is ultimately generated. Specifically, we analyze the top-$k$ tokens ranked by logit scores in cases where the model initially outputs “unsure.” We then apply a filtering procedure to remove uninformative candidates (e.g., “unsure”, null strings, or stop words), and identify whether the remaining tokens contain the correct answer.
This procedure allows us to estimate the gap between knowledge storage and expression, revealing how often the model possesses the correct information but fails to surface it due to confidence calibration or decoding dynamics. The details of this method are described below.
To quantify this phenomenon, we propose a simple two-stage decoding procedure that filters out “unsure”-related tokens and re-invokes the model for answer generation. The decoding pipeline is outlined in Algorithm 1.